OpenStudy (anonymous):

how does one determine the displacement of a graph? how does one determine the displacement of a graph? @Physics

OpenStudy (asnaseer):

can you please give a bit more context to your question.

OpenStudy (anonymous):

if there is a velocity time graph the take are under the graph(integration) if there is acceleration time graph then u will get velocity then u can apply newtons law acc. to the ques.

OpenStudy (turingtest):

@heena No, integrating a velocity function over time will give distance (scalar), not displacement (vector). To find displacement from a graph all that is needed is to subtract the initial position from the final one.

OpenStudy (aravindg):

@heena u free??

OpenStudy (anonymous):

okhezz.. distance is the magnitude of diplacement rit.. i done mistake sorry

OpenStudy (amistre64):

y = (x+3)^2 is the graph of y=x^2 displaced by 3 units to the left ....

OpenStudy (turingtest):

 in this graph displacement would be found by subtracting \[x(t_2)-x(t_1)=x_2-x_1\]